OpenStudy (anonymous):

Find the values of the trigonometric functions of t from the given information. sec(t) = 2, sin(t) < 0 sin(t) and tan(t)

OpenStudy (anonymous):

\[\sec(t)=2\implies \cos(t)=\frac{1}{2}\] for starters

OpenStudy (anonymous):

you need \(\sin(t)\) which you get either by remembering that where \(\cos(t)=\frac{1}{2},\sin(t)=\pm\frac{\sqrt{3}}{2}\) or by pythagoras

OpenStudy (anonymous):



OpenStudy (anonymous):

and you get \(\tan(t)\) via \(\tan(t)=\frac{\sin(t)}{\cos(t)}\)

OpenStudy (anonymous):

Thanks